# Identifying multi-compartment Hodgkin-Huxley models with high-density extracellular voltage recordings

### v1

@@ Appendix B Additional Details and Figures for Section 4
The extracellular resistivity was assumed to be homogeneous and equal to $300\,\Omega\cdot\mathrm{cm}$, and the specific membrane capacitance was fixed to $C_{m}=1\,\mu\mathrm{F}/\mathrm{cm}^{2}$ across all compartments. While we assume homogeneous extracellular resistivity, this simplification neglects known spatial variability [Gold et al., 2006], which could be addressed in future work. Extracellular voltages were recorded at a sampling rate of 40 kHz.
All voltages and channel states were initialized at their steady-state values at rest, and thus we assumed the initial state to be known in the EKF. We set the initial covariance matrix of the EKF to $\Sigma_{0}=0.1I$. Observation noise was also assumed to be known, as commercial devices that record extracellular potentials typically provide noise specifications [Ye et al., 2024, Multi Channel Systems MCS GmbH, 2020].
@@
B.1 Validation on Single-Compartment Neuron Model
We evaluated a single-compartment neuron model as a baseline for benchmarking. The model consists of a cylindrical compartment of length $24\,\mu\mathrm{m}$ and radius $12\,\mu\mathrm{m}$, equipped with classical Hodgkin–Huxley (HH) sodium, potassium, and leak channels [Hodgkin and Huxley, 1952]. The maximum conductances and reversal potentials are provided in Table 1. A constant current of amplitude $2.5\,\mathrm{nA}$ was injected for 20 ms.
@@ B.1 Validation on Single-Compartment Neuron Model
Reversal potential (mV) 53.0 $-107.0$ $-88.5188$
Table 1: Biophysical parameters of the single-compartment neuron model.
@@
As the neuron–electrode distance increased, the signal-to-noise ratio (SNR) decreased, resulting in higher bias and variance in inferred conductances (fig. B.1B). At short distances with high SNR, both MSE minimization and EKF yielded comparable estimates, indicating that EKF’s linearization introduced minimal error (fig. B.1C). As observation noise increased, estimation accuracy declined (fig. B.1A), but even at $30\,\mu\mathrm{V}$ RMS—a challenging noise regime (fig. B.2C)—estimates remained close to the ground truth. Notably, the true voltage remained within two standard deviations of the mean of the reconstructed voltage traces for most of the signal (fig. B.3), indicating robustness even under challenging conditions.
Figure B.1: Validation on synthetic single-compartment. A: Distributions of inferred maximum conductances under different observation noise levels $\sigma_{\text{obs}}$ (colored dashed lines indicate the means). B: Distributions of inferred maximum conductances for different cell-probe distances (colored dashed lines indicate the means). C: Distributions of inferred maximum conductances for MSE minimization and EKF, with $\sigma_{\text{obs}}=3\,\mu\text{V}$ and $d=10\mu\mathrm{m}$.
@@
Figure B.2: Extracellular voltage traces simulated for different noise levels in the single-compartment model. A: $\sigma_{obs}=3\mu\mathrm{V}$, B: $\sigma_{obs}=10\mu\mathrm{V}$, C: $\sigma_{obs}=30\mu\mathrm{V}$ represents a challenging noise regime. The electrode is $10\mu\mathrm{m}$ away from the cell.
@@
Figure B.3: Recovery of membrane voltage for different observation noise levels in the single-compartment model. The mean and standard deviation are estimated by considering 40 datasets simulated at fixed observation noise. Even in challenging noise regimes ($\sigma_{obs}=30\mu\mathrm{V}$), the true voltage remains within two standard deviations of the mean of the reconstructed voltage traces for most of the signal.
B.2 Validation on Multi-Compartment Neuron Model
@@
We also evaluated a multi-branch neuron model to test conductance and geometry inference. The model consists of three branches, each discretized into four cylindrical compartments of radius $2\,\mu\mathrm{m}$, for a total of 12 compartments. Each compartment contains classical Hodgkin-Huxley sodium, potassium, and leak channels [Hodgkin and Huxley, 1952]. Reversal potentials are shared across all branches (Table 2), while conductances and lengths vary by branch (Table 3). A current of $0.5\,\mathrm{nA}$ was injected in the first compartment of branch 1 (soma) for 20 ms.
@@
All the maximum conductances were inferred accurately for all $40$ datasets simulated with $\sigma_{obs}=1\mu\mathrm{V}$ (fig. B.4).
@@
Table 3: Maximum conductances (in S/cm2) and branch lengths (in $\mu$m) in the multi-branch neuron model.
Figure B.4: Validation on synthetic multi-branch neuron. Inferred maximum conductances $g_{\mathrm{L}}$, ${g}_{\mathrm{K}}$ and ${g}_{\mathrm{Na}}$ for all the branches, across 40 datasets simulated with $\sigma_{\text{obs}}=1\,\mu\text{V}$.
B.3 Robustness to Model Misspecification
@@
To test robustness to missing channels, we simulated data from a single-compartment model (length $10\mu\mathrm{m}$, radius $1\mu\mathrm{m}$) containing M-type potassium (KM) and L-type calcium (CaL) channels (Table 5) in addition to Hodgkin-Huxley (HH) channels (Table 4). Inference was performed using a model containing only the HH channels. A step current of $0.004\,\mathrm{nA}$ was injected for 20 ms. Extracellular voltages were recorded at a single electrode located $10\,\mu\mathrm{m}$ away from the soma, with a noise level of $\sigma_{obs}=0.1\,\mu\mathrm{V}$. Biophysical parameters for the simulation model are shown in Table 6.
@@
For MSE-based optimization, we used Adam and reported the result with the lowest final MSE.
@@
Table 4: Rate functions for sodium (Na+) and potassium (K+) channels used in the single-compartment model for the missing channel experiment.
@@
Table 5: Rate functions for M-type potassium (KM) and L-type calcium (CaL) channels used in the single-compartment model for the missing channel experiment.
@@
Reversal potential (mV) 50.0 $-90.0$ $-70.0$ $-90.0$ $120$
Table 6: Biophysical parameters for the single-compartment neuron model used in the missing channel experiment.
@@
To evaluate robustness to incorrect kinetics, we generated data using gating functions from a retinal ganglion cell (RGC) model (Table 8) and performed inference using the original Hodgkin-Huxley kinetics [Hodgkin and Huxley, 1952]. The neuron was a single cylindrical cable (length $24\,\mu\mathrm{m}$, radius $2\,\mu\mathrm{m}$), discretized into four compartments. A constant current of $0.5\,\mathrm{nA}$ was injected into the first compartment (soma) for 20 ms. Extracellular voltages were recorded at 10 different locations with noise level $\sigma_{obs}=0.1\,\mu\mathrm{V}$. Biophysical parameters for the simulation model are shown in Table 7.
@@
The EKF smoothed posterior mean of the voltage recovers the true voltage trace in each of the four compartments, whereas MSE minimization fails (fig. B.5).
@@
Figure B.5: Robustness to misspecified channel kinetics. The EKF smoothed posterior mean of the voltage recovers the true voltage trace in each of the four compartments, whereas MSE minimization fails.
Appendix C Unidentifiability of 3D compartment locations
Our multi-compartment model represents the neuron as a collection of compartments, each treated as a point source of transmembrane current located at its center. Hence, estimating a neuron’s spatial configuration reduces to estimating the 3D positions of its compartments. According to Eq. 3, the extracellular potential measured at a given electrode depends only on the transmembrane currents and the inverse distance to each compartment. Therefore, if the distances between each compartment and all electrodes remain unchanged, the resulting extracellular signals will be identical.
We first consider the case where all electrodes lie on a single line (fig. C.1A). Fix a single compartment at some position in 3D space (black). The set of all locations that preserve its distance to a given site is a sphere. The intersection of these spheres—corresponding to multiple sites aligned on a line—is a circle. Thus, any position on this circle (blue dashed circle) yields the same set of distances (and hence the same extracellular potentials), rendering the compartment’s position unidentifiable up to a rotation around the axis defined by the recording sites.
Next, suppose the electrodes lie on two parallel lines, forming a planar grid (fig. C.1B). In this case, for a given compartment (black), the set of equidistant locations from the first line of sites lies on a circle (blue), and similarly for the second line (green). The only positions that simultaneously preserve distances to both sets of sites lie at the intersection of these two circles. This intersection consists of exactly two points: the true location of the compartment and its mirror reflection across the plane defined by the electrodes (grey).
This ambiguity persists even with more than two lines of electrodes, as long as they all lie in the same plane. For example, fig. C.1C shows three parallel lines of recording sites in a planar grid, viewed from above. The mirror-reflected location remains the only position (besides the true one) that produces identical distances to all recording sites.
Figure C.1: Unidentifiability of the positions. A: The distance between the compartment (black) and each electrode (blue) is unchanged as it moves along the blue dashed circle. It follows that any position of the compartment on the circle yields the same observations at all the electrodes. B: Compartment’s distance to each blue (resp. green) electrode is unchanged as it moves along the blue (resp. green) dashed circle. Hence, the grey location is the only other position that yields the same observations as the black one at all electrodes. C: Panel B viewed from above with now three parallel lines containing electrodes (green, blue, orange ones). Since all the electrodes lie in a 2D-plane $(z=0)$, the only position that yields the same observations as the true location at each electrode is the true location’s reflection with respect to the plane $z=0$.
Appendix D Retinal Ganglion Cell (RGC) Model
We use a morphologically detailed model of a retinal ganglion cell (RGC), based on a reconstruction from Vilkhu et al. [2024]. The morphology was originally provided as a NEURON .hoc file, [^1] https://github.com/ramanvilkhu/rgc_simulation_multielectrode which we converted to SWC format for compatibility with Jaxley. The model includes the full soma and dendritic arbor, and was discretized into 122 compartments, assigning one compartment per branch.
The ion channel model consists of Hodgkin–Huxley (HH) conductances: fast sodium, delayed-rectifier potassium, and leak channels. Voltage-dependent rate constants are listed in Table 8, and channel reversal potentials in Table 9. Region-specific maximum conductances for soma and dendrites are given in Table 10. All values were taken directly from Vilkhu et al. [2024], with a minor adjustment to the leak reversal potential to ensure resting membrane potentials near $-65$ mV across compartments. Note that Vilkhu et al. [2024] also consider two additional channels (a calcium channel and a calcium-gated potassium channel), which we neglected in this experiment. Including these channels could be the subject of future work.
@@ Appendix D Retinal Ganglion Cell (RGC) Model
To generate synthetic extracellular recordings, we applied a constant somatic current injection of $1.0$ nA for 10 ms. Voltage recordings were simulated using a synthetic planar multi-electrode array (MEA) with a sampling rate of 40 kHz.
We fit this model using our diagEKF implementation. All voltages and channel states were initialized at their steady-state values at rest, and thus we assumed the initial state to be known in diagEKF. We set the initial covariance matrix to $\Sigma_{0}=0.1I$. Observation noise was also assumed to be known, as commercial devices that record extracellular potentials typically provide noise specifications [Ye et al., 2024, Multi Channel Systems MCS GmbH, 2020]. Since the state dynamics were well-specified and deterministic, we fixed the standard deviation of the membrane potentials and channel states to small values: $\sigma_{v}=0.001\,\mathrm{mV}/\mathrm{ms}^{1/2}$, and $\sigma_{\text{state}}=0.0001\ \mathrm{ms}^{-1/2}$ (same for all the channels) for all compartments. These values ensured numerical stability while avoiding degenerate gradients during optimization.
We optimized conductance parameters using the Adam optimizer. We tried multiple learning rates and initialized values uniformly at random within the bounds shown in Table 11. We ran 100 random initializations with different seeds and selected the result that achieved the highest marginal log-likelihood.
All maximum conductances were accurately inferred across all 10 datasets simulated at a fixed noise level of $\sigma_{\text{obs}}=0.1\,\mu\mathrm{V}$, with only minor biases observed for a few conductances (fig. D.1). As expected, increasing the observation noise to $\sigma_{\text{obs}}=10\,\mu\mathrm{V}$ generally led to higher bias and variance of the estimates, though the inferred values remained close to the ground truth (fig. D.2). In both noise conditions, membrane voltage was accurately recovered throughout the cell (fig. D.3), and the marginal log-likelihood converged to its optimal value (fig. D.4).
@@
Figure D.1: Inference of retinal ganglion cell (RGC) properties given multi-electrode array (MEA) recordings. Estimation of all the maximum conductances for 10 optimization runs under noise level $\sigma_{\text{obs}}=0.1\mu\text{V}$. All the maximum conductances were recovered accurately, with minor bias occasionally.
Figure D.2: Impact of the observation noise on inference for the RGC model. Distributions of inferred maximum conductances for $\sigma_{\text{obs}}\in\{0.1,10\}\,\mu\text{V}$ (colored dashed lines indicate the means). Increasing the observation noise generally led to higher bias and variance of the estimates, though the inferred values remained close to the ground truth.
Figure D.3: Membrane voltage recovery for the RGC model in six compartments. A: soma, B: compartment 31, C: compartment 61, D: compartment 91, E: compartment 106, F: compartment 122. The ordering of the compartments is determined by the SWC file. For both low and higher noise levels, the membrane voltage was recovered accurately across the cell.
Figure D.4: Marginal log-likelihood trace during optimization across 10 runs for the RGC model. A: $\sigma_{obs}=0.1\mu\mathrm{V}$, B: $\sigma_{obs}=10\mu\mathrm{V}$. The marginal log-likelihood converged to its optimal value for both noise regimes.
Appendix E CA1 Pyramidal Neuron (C10 Model)
We use the C10 model of a CA1 pyramidal neuron, originally developed by Cutsuridis et al. [2010]. The morphology is provided as a NEURON .hoc file, [^2] https://github.com/tomko-neuron/HippoUnit/tree/master/Cutsuridis_CA1_2010, which we converted to SWC format for compatibility with Jaxley. This file specifies the full 3D morphology, including the diameters and lengths of all the branches (Table 14). As shown in fig. E.1, the neuron is divided into 15 anatomical sections.
Following Cutsuridis et al. [2010], we applied the compartmentalization heuristic of Carnevale and Hines [2006] to determine the number of compartments per section. Specifically, a section of length $L$ is discretized into

### v2

@@ Appendix E Additional Details and Figures for Section 4
The extracellular resistivity was assumed to be homogeneous and equal to $300\,\Omega\cdot\mathrm{cm}$, and the specific membrane capacitance was fixed to $C_{m}=1\,\mu\mathrm{F}/\mathrm{cm}^{2}$ across all compartments. While we assume homogeneous extracellular resistivity, this simplification neglects known spatial variability [gold2006origin], which could be addressed in future work. Extracellular voltages were recorded at a sampling rate of 40 kHz.
All voltages and channel states were initialized at their steady-state values at rest, and thus we assumed the initial state to be known in the EKF. We set the initial covariance matrix of the EKF to $\mathbf{\Sigma}_{0}=0.1\mathbf{I}$. Observation noise was also assumed to be known, as commercial devices that record extracellular potentials typically provide noise specifications [ye2024ultra, MEA2100manual].
@@
Figure E.1: Validation on synthetic single-branch. A: Distributions of inferred maximum conductances under different observation noise levels $\sigma_{\text{obs}}$ (colored dashed lines indicate the means). B: Distributions of inferred maximum conductances for different cell-probe distances (colored dashed lines indicate the means). C: Distributions of inferred maximum conductances for various inference methods with $\sigma_{\text{obs}}=1\,\mu\text{V}$ and $d=5\mu\mathrm{m}$ (orange lines indicate the means).
Figure E.2: Recovery of membrane voltage with diagonal EKF in the single-branch model for compartments 1 (left), 5 (center) and 10 (right). The mean and standard deviation are estimated by considering 40 randomly simulated datasets at fixed observation noise $\sigma_{\text{obs}}=1\mu\mathrm{V}$. Despite a small bias in the estimation of the parameters, the voltage traces are still reconstructed accurately.
E.1 Validation on single-branch model
We evaluated a single-branch cell model as a baseline for benchmarking. This branch consists of 10 cylindrical compartments of length $24\,\mu\mathrm{m}$ and radius $2\,\mu\mathrm{m}$, equipped with standard Hodgkin–Huxley (HH) sodium, potassium, and leak channels [hodgkin1952quantitative]. The maximum conductances and reversal potentials are provided in Table 1. A constant current of amplitude $1.5\,\mathrm{nA}$ was injected for 20 ms.
@@ E.1 Validation on single-branch model
Reversal potential (mV) 53.0 107.0-107.0 88.5188-88.5188
Table 1: Biophysical parameters of the single-branch model.
@@
As the neuron–electrode distance increased, the signal-to-noise ratio (SNR) decreased, resulting in higher bias and variance in inferred conductances (fig.˜E.1B). At short distances with high SNR, we observed that MSE minimization, standard EKF and block-diagonal EKF yielded comparable estimates, indicating that EKF’s linearization introduced minimal error and confirming the validity of the block-diagonal approximation (fig.˜E.1C). The diagonal EKF introduces a slightly larger bias, but its voltage reconstructions remain highly accurate (fig.˜E.2). This is consistent with our analysis in Appendix B: only the first (soma) compartment lies near the electrode, with all others much farther away, so cross-visibility is minimal and a diagonal covariance approximation is sufficient. As observation noise increased, estimation accuracy declined (fig.˜E.1A), but even at $30\,\mu\mathrm{V}$ RMS—a challenging noise regime (fig.˜E.3C)—estimates remained close to the ground truth. Notably, the induced variance on the reconstructed voltage traces is small (fig.˜E.4), indicating robustness even under challenging conditions.
@@
Figure E.3: Extracellular voltage traces simulated for different noise levels in the single-branch model. A: $\sigma_{obs}=1\mu\mathrm{V}$, B: $\sigma_{obs}=10\mu\mathrm{V}$, C: $\sigma_{obs}=30\mu\mathrm{V}$ represents a challenging noise regime. The electrode is $5\mu\mathrm{m}$ away from the cell.
@@
Figure E.4: Recovery of membrane voltage for different observation noise levels in the single-branch model. The mean and standard deviation are estimated by considering 40 datasets simulated at fixed observation noise. Even in challenging noise regimes ($\sigma_{obs}=30\mu\mathrm{V}$), the induced variance on the voltage trace reconstruction is negligible.
E.2 Validation on multi-branch neuron model
@@
We also evaluated a multi-branch neuron model to test conductance and geometry inference. The model consists of three branches, each discretized into four cylindrical compartments of radius $2\,\mu\mathrm{m}$, for a total of 12 compartments. Each compartment contains standard Hodgkin-Huxley sodium, potassium, and leak channels [hodgkin1952quantitative]. Reversal potentials are shared across all branches (Table 2), while conductances and lengths vary by branch (Table 3). The axial resistivity was set to $150\ \Omega\cdot\mathrm{cm}$. A current of $2\,\mathrm{nA}$ was injected in the first compartment of branch 1 (soma) for 20 ms.
@@
With EKF, all the maximum conductances were inferred accurately for all $40$ datasets simulated with $\sigma_{obs}=1\mu\mathrm{V}$ (fig.˜E.5).
With the diagonal EKF, the recovered conductances exhibit a more pronounced bias than in the single-branch case, particularly in branches 2 and 3 (fig.˜E.6), although the voltage traces remain accurate up to small shifts in these branches (fig.˜E.7), and the compartment center locations were inferred correctly (fig.˜E.8). In this multi-branch configuration, the diagonal approximation retains decent accuracy because, despite the relatively dense electrode layout, the long-compartment geometry reduces instances where electrodes are simultaneously close to distinct compartments. Since each compartment only has two neighbors (except at the branching node), this limits off-diagonal measurement coupling. As detailed in Appendix B, this geometry ensures that most of the off-diagonal entries of $(\mathbf{H}^{v})^{\top}\mathbf{H}^{v}$ remain significantly smaller than terms near the diagonal (fig.˜E.9). Consequently, the diagonal approximation introduces some bias but continues to provide useful and stable inference.
@@
[…]
Figure E.5: Validation of the EKF on a synthetic multi-branch neuron. The maximum conductances $g_{\mathrm{L}}$, ${g}_{\mathrm{K}}$ and ${g}_{\mathrm{Na}}$ were inferred accurately for all the branches, across 40 datasets simulated with $\sigma_{\text{obs}}=1\,\mu\text{V}$.
Figure E.6: Inferred parameters using the diagonal EKF on a synthetic multi-branch neuron. We show results across 40 datasets simulated with $\sigma_{\text{obs}}=1\,\mu\text{V}$. The recovered conductances exhibit a more pronounced bias than in the single-branch case, particularly in branches 2 and 3.
Figure E.7: Reconstructed voltage for first (top) and last (bottom) compartment of branch 1 (left), 2 (middle) and 3 (right) using the diagonal EKF on a synthetic multi-branch neuron. Voltage traces are averaged over 40 randomly simulated datasets with $\sigma_{\text{obs}}=1\mu\text{V}$. Although conductance estimates exhibit noticeable bias, the reconstructed voltages remain accurate up to small shifts, consistent with a regime in which the diagonal approximation remains adequate.
Figure E.8: Diagonal EKF accurately recovers compartment center locations. Estimated 3D positions are averaged over 40 randomly simulated datasets with $\sigma_{\text{obs}}=1\mu\text{V}$. Although conductance estimates exhibit noticeable bias, the reconstructed compartment locations remain accurate, consistent with a regime in which the diagonal approximation remains adequate.
Figure E.9: Heatmap (in log scale) of $\mathbf{H}^{v\top}\mathbf{H}^{v}$, the voltage block of $\mathbf{H}^{\top}\mathbf{H}$. The largest coefficients occur between adjacent compartments and at branch points, yielding an approximately tridiagonal structure. Off-diagonal entries corresponding to distant compartments are substantially smaller, reflecting the long-compartment morphology, where few electrodes are simultaneously close to two non-neighboring compartments. This structure helps explain why the diagonal approximation remains reasonably effective in this setting.
E.3 Robustness to Model Misspecification
@@
To test robustness to missing channels, we simulated data from a four-compartment branch model (total length $96\mu\mathrm{m}$, radius $2\mu\mathrm{m}$) containing M-type potassium (KM) and L-type calcium (CaL) channels (Table 5) in addition to Hodgkin-Huxley (HH) channels (Table 4). Inference was performed using a model containing only the HH channels. A constant current was injected for 20 ms. Extracellular voltages were recorded at ten electrodes with a noise level of $\sigma_{obs}=0.1\,\mu\mathrm{V}$. Biophysical parameters for the simulation model are shown in Table 6.
@@
For MSE-based optimization, we used Adam [kingma2014adam] and reported the result with the lowest final MSE.
@@
Table 4: Rate functions for sodium (Na+) and potassium (K+) channels used in the single-branch model for the missing channel experiment.
@@
Table 5: Rate functions for M-type potassium (KM) and L-type calcium (CaL) channels used in the single-branch model for the missing channel experiment.
@@
Reversal potential (mV) 50.0 90.0-90.0 70.0-70.0 90.0-90.0 $120$
Table 6: Biophysical parameters for the single-branch neuron model used in the missing channel experiment.
@@
To evaluate robustness to incorrect kinetics, we generated data using gating functions from a retinal ganglion cell (RGC) model (Table 8) and performed inference using the original Hodgkin-Huxley kinetics [hodgkin1952quantitative]. The neuron was a single cylindrical cable (length $24\,\mu\mathrm{m}$, radius $2\,\mu\mathrm{m}$), discretized into four compartments. A constant current was injected into the first compartment (soma) for 20 ms. Extracellular voltages were recorded at 10 different locations with noise level $\sigma_{obs}=0.1\,\mu\mathrm{V}$. Biophysical parameters for the simulation model are shown in Table 7.
@@
The EKF smoothed posterior mean of the voltage recovers the true voltage trace in each of the four compartments, whereas MSE minimization fails (fig.˜E.10).
@@
Figure E.10: Robustness to misspecified channel kinetics. The EKF smoothed posterior mean of the voltage recovers the true voltage trace in each of the four compartments, whereas MSE minimization fails.
Appendix F Retinal Ganglion Cell (RGC) Model
We use a morphologically detailed model of a retinal ganglion cell (RGC), based on a reconstruction from vilkhu2024understanding. The morphology was originally provided as a NEURON .hoc file, [^1] https://github.com/ramanvilkhu/rgc_simulation_multielectrode which we converted to SWC format for compatibility with Jaxley. The model includes the full soma and dendritic arbor, and was discretized into 122 compartments, assigning one compartment per branch.
The ion channel model consists of Hodgkin–Huxley (HH) conductances: fast sodium, delayed-rectifier potassium, and leak channels. Voltage-dependent rate constants are listed in Table 8, and channel reversal potentials in Table 9. Region-specific maximum conductances for soma and dendrites are given in Table 10. All values were taken directly from vilkhu2024understanding, with a minor adjustment to the leak reversal potential to ensure resting membrane potentials near 65-65 mV across compartments. Note that vilkhu2024understanding also consider two additional channels (a calcium channel and a calcium-gated potassium channel), which we neglected in this experiment. Including these channels could be the subject of future work.
@@ Appendix F Retinal Ganglion Cell (RGC) Model
To generate synthetic extracellular recordings, we applied a constant somatic current injection for 10 ms. Voltage recordings were simulated using a synthetic planar multi-electrode array (MEA) with a sampling rate of 40 kHz.
We fit this model using our diagEKF and block-diagEKF implementations. All voltages and channel states were initialized at their steady-state values at rest, and thus we assumed the initial state to be known. We set the initial covariance matrix to $\mathbf{\Sigma}_{0}=0.1\mathbf{I}$. Observation noise was also assumed to be known, as commercial devices that record extracellular potentials typically provide noise specifications [ye2024ultra, MEA2100manual]. Since the state dynamics were well-specified and deterministic, we fixed the standard deviation of the membrane potentials and channel states to small values: $\sigma_{v}=0.001\,\mathrm{mV}/\mathrm{ms}^{1/2}$, and $\sigma_{\text{state}}=0.0001\ \mathrm{ms}^{-1/2}$ (same for all the channels) for all compartments. These values ensured numerical stability while avoiding degenerate gradients during optimization.
We optimized conductance parameters using the Adam optimizer. We tried multiple learning rates and initialized values uniformly at random within the bounds shown in Table 11. We ran over 100 random initializations with different seeds and selected the result that achieved the highest marginal log-likelihood.
With the block-diagonal EKF, all maximum conductances were accurately inferred (F.1) and the voltage traces faithfully reconstructed across the cell (F.2).
@@
Figure F.1: Inference of retinal ganglion cell (RGC) properties given multi-electrode array (MEA) recordings (block-diagonal EKF). Estimation of all the maximum conductances at noise level $\sigma_{\text{obs}}=0.1\mu\text{V}.$
Figure F.2: Membrane voltage recovery for the RGC model in six compartments (bloc-diagonal EKF). A: soma, B: compartment 31, C: compartment 61, D: compartment 91, E: compartment 106, F: compartment 122. The ordering of the compartments is determined by the SWC file.
Appendix G CA1 Pyramidal Neuron (C10 Model)
We use the C10 model of a CA1 pyramidal neuron, originally developed by cutsuridis2010encoding. The morphology is provided as a NEURON .hoc file, [^2] https://github.com/tomko-neuron/HippoUnit/tree/master/Cutsuridis_CA1_2010, which we converted to SWC format for compatibility with Jaxley. This file specifies the full 3D morphology, including the diameters and lengths of all the branches (Table 14). As shown in fig.˜G.1, the neuron is divided into 15 anatomical sections.
Following cutsuridis2010encoding, we applied the compartmentalization heuristic of carnevale2006neuron to determine the number of compartments per section. Specifically, a section of length $L$ is discretized into